# inkscapelatex=false

### v2

@@ 6.3 Actor Critic
Advantage estimates $A_{t}$ are computed from multi-step discounted returns using generalized advantage estimation (GAE; [Schulman et al., 2015]) with parameters $\gamma$=0.99 and $\lambda$=0.95. Training is performed in the asynchronous actor–learner framework of Sample Factory [Petrenko et al., 2020].
6.4 Training configuration
We trained all agents using the Sample Factory framework [Petrenko et al., 2020] with an asynchronous distributed setup. Each experiment used 32 workers running 8 environments per worker, with experience decorrelated for up to 120 seconds. The environment applied a frame-skip of 8, i.e., each selected action was repeated for 8 frames before the next observation was retrieved.
We used the Asynchronous Proximal Policy Optimization (APPO) algorithm [Petrenko et al., 2020] with discount factor $\gamma=0.99$ and GAE parameter $\lambda=0.95$. Training was performed with a rollout length of 64, a recurrence length of 64, a batch size of 2048, and two batches per epoch. The optimizer used a learning rate of $2\times 10^{-4}$.
The agent architecture followed the design summarized in Table A3. The visual encoder was a ResNet backbone pretrained with an LSTM core on DMLab. For the control experiment, we additionally tested the second convolutional block of ResNet-18 pretrained on ImageNet [He et al., 2015]. The dentate gyrus (DG) module applied batch normalization followed by ReLU activation with an intercept of 2.43, producing 16 sparse feature channels. The decoder consisted of two fully connected layers of 128 units each with ReLU activations.
We employed Population-Based Training (PBT) with four concurrent policies. Policy replacement occurred when the performance gap exceeded 0.05 in relative terms or 0.2 in absolute reward units. Mutations began after 10 million environment steps and were evaluated every 2 million steps. A policy lag tolerance of 35 was used to maintain stability.
Training ran for 108,000 seconds, with checkpoints saved every 120 seconds and milestone checkpoints every 5400 seconds. All experiments were run on CPU, with affinity pinning disabled and input normalization turned off. We used random seeds [1111, 2222, 3333, 4444, 5555]. All other hyperparameters followed the default settings provided in Sample Factory for the DeepMind Lab environment.
6.5 Skaggs’ Spatial Information Measure
We quantify how much information a unit’s activity conveys about the agent’s location in our discrete‐time simulation by a “bits per step” version of Skaggs’ spatial information Skaggs et al. [1992]. Let the environment be divided into $N$ spatial bins, and define:
@@ 6.5 Skaggs’ Spatial Information Measure
7 Supplementary Figures
@@
Figure A1: Robustness to changes of visual encoder. Training performance of CA3 agent with the ResNet layer 2 output as visual encoder. Random seeds: [1111, 2222, 3333, 4444, 5555].
@@
Figure A2: Robustness to changes of realization of environment. Training performance on 5 other randomly generated maps. Lines with the same color are results from random seeds [0, 42]
@@
Figure A3: Comparison to state-space model (SSM) based architectures. Training performance of agents with fixed SSM cores described by Gu et al. [2020]. legt: Legendre bases with fixed memory horizon. legs: Legendre bases with infinite memory from the beginning of an episode. lagt: Laguerre bases, i.e. memory with decay. All three SSMs were implemented with the same state size as our CA3 module, each input feature is expanded into 64+8-1=71 states. The SSMs were fixed, with three additional learnable parameter controlling the input, recurrent scales and timestep length for discretizations, as a common practice described by Gu et al. [2021].
@@
Figure A4: More examples of rate maps from decoder layer 1 in the last training epoch. Showing the random 16 units out of the 50 units with the best SI. Unit id displayed on top.
@@
Figure A5: More examples of rate maps from decoder layer 2 in the last training epoch. Showing the random 16 units out of the 50 units with the best SI. Unit id displayed on top.
@@
Figure A6: Distributions of spatial information (SI) rates becomes heavy tailed over learning. SI rate distributions for different layers (brain regions rows) over the progress of learning (colors) and change of reward site (purple). Y axes: unit fraction density.
@@
Figure A7: Population representations stabilize over learning and remap. PV correlations across training. Although performance saturates by epochs 5–6, representational changes continue. The last epoch shows the effect of learning a new goal location.
Figure A8: Vision centered behavioral strategy in LSTM-based agents. Occupancy map and mean movement direction (white) of LSTM-based agent’s trajectories across 5 epochs of training.
@@
Figure A9: Example ratemap of LSTM output units. Place fields from the LSTM core (randomly selected units). Each panel shows the normalized spatial activity of one unit.
@@
Figure A10: LSTM-based agent exhibits less input orthogonalization. Population vector correlation of LSTM inputs across epochs. Color indicates Pearson correlation; rows show epochs 0–4.
@@
Figure A11: LSTM-based agent exhibits less input sparseness. Learned dense input features in LSTM agent. Each tile shows the spatial receptive field of one feature.
Figure A12: No clear distance-dependent representation of LSTM-based agents. A Displacement-dependent Pearson correlation kernel of population vectors. Axes show spatial shifts $(\Delta x,\Delta y)$; values are averaged correlations. Rows correspond to different layers (brain regions) of the network, columns to learning stages (epochs). B Left: summary of population vector correlation as a function of distance. Right: summary of population vector correlation as a function of orientation. Radial coordinate range: (-0.5,1.1).
@@
Figure A13: Robustness to changes in training parameters. Training performance with different environment action repeats and number of policies in population based training. Left: environment frame skip / action repeat = 8. Right: environment frame skip / action repeat = 4. The frame skip controls the fine-graininess of actions, which is not required for the current navigation task. On the other hand, larger frame skip effectively make sequences propagate over longer traversals.

### v3

@@ B.4 Actor Critic
Here $R_{t}$ and the advantage estimates $A_{t}$ are computed using V-trace returns (Espeholt et al., 2018), which correct for off-policy updates arising in the asynchronous actor–learner setup of Sample Factory (Petrenko et al., 2020). This formulation—known as Asynchronous PPO (APPO)—combines PPO’s clipped surrogate objective (Schulman et al., 2017) with IMPALA’s V-trace corrections, enabling scalable training with many actors while maintaining stable policy updates.
B.5 Training configuration
We trained agents using sample-factory (Petrenko et al., 2020) with the following setup.
Environment.
frameskip 8, repeating action for 8 frames until getting the next observation. Each run used 32 workers $\times$ 8 envs/worker, with decorrelation up to 120s.
Algorithm.
APPO (Espeholt et al., 2018), $\gamma=0.99$, rollout length 64, recurrence 64, batch size 2048, 2 batches/epoch. Optimizer learning rate $2\!\times\!10^{-4}$.
Architecture.
Visual encoder: pretrained ResNet on DMLab and the second layer of ResNet pretrained on ImageNet DG: batchnorm + ReLU, intercept 2.43, with 16 features. Decoder: 2 MLP layers of size 128.
Population Based Training.
Enabled PBT with 4 policies, replacement gaps 0.05 (relative) and 0.2 (absolute), mutation start after 10M steps, period every 2M steps. Policy lag tolerance set to 35.
Logging and checkpoints.
Training for 108k seconds, milestones every 5400s.
Miscellaneous.
Seeds: [1111,2222,3333,4444,5555]. Device: CPU. Affinity pinning disabled. Inputs not normalized. Other parameters were default in sample-factory for DeepMind Lab.
B.6 Skaggs’ Spatial Information Measure
We quantify how much information a unit’s activity conveys about the agent’s location in our discrete‐time simulation by a “bits per step” version of Skaggs’ spatial information (Skaggs et al., 1992). Let the environment be divided into $N$ spatial bins, and define:
@@ B.6 Skaggs’ Spatial Information Measure
Appendix C Supplementary Table
Table A3: The effect of parameter R at different running speed (frameskip). Performance is measured as the success rate at 150 million training frames. This is a summary of Fig. A1.
Frameskip R=1 R=4 R=8 R=12 R=16 R=32
[…]
Table A4: The performance at different noise level. Performance is measured as the success rate at 100 million training frames. This is a summary of Fig. A5.
Model $\backslash$ Noise 0 10 20 40 80
DG+CA3 0.693 $\pm$ 0.137 0.492 $\pm$ 0.167 0.539 $\pm$ 0.108 0.685 $\pm$ 0.186 0.477 $\pm$ 0.072
Dense LSTM 0.638 $\pm$ 0.063 0.707 $\pm$ 0.178 0.495 $\pm$ 0.010 0.449 $\pm$ 0.064 0.373 $\pm$ 0.008
Appendix D Supplementary Figures
Figure A1: Effect of parameter R with different running speed. Training performance of CA3 agent with L=64 different R. Random seeds: [$11$, $22$, $33$, $44$, $55$].
@@ Appendix D Supplementary Figures
Figure A2: Robustness of visual encoder. Training performance of CA3 agent with the ResNet layer 2 output as visual encoder. Random seeds: [$11$, $22$, $33$, $44$, $55$].
@@
Figure A3: Training performance on 5 other randomly generated maps. Lines with the same color are results from random seeds [0, 42]
@@
Figure A4: Training performance of agents with fixed SSM cores described by Gu et al. (2020). legt: Legendre bases with fixed memory horizon. legs: Legendre bases with infinite memory from the beginning of an episode. lagt: Laguerre bases, i.e. memory with decay. All three SSMs were implemented with the same state size as our CA3 module, each input feature is expanded into 64+8-1=71 states. The SSMs were fixed, with three additional learnable parameter controlling the input, recurrent scales and timestep length for discretizations, as a common practice described by Gu et al. (2021).
Figure A5: Training performance of agents under different noise levels. Independent Gaussian noise of different sigmas was added to the input image pixels (intensity 0-255). Random seeds: [$11$, $22$, $33$, $44$, $55$].
@@
Figure A6: Rate maps from decoder layer 1 in the last training epoch. Showing the random 16 units out of the 50 units with the best SI. Unit id displayed on top.
@@
Figure A7: Rate maps from decoder layer 2 in the last training epoch. Showing the random 16 units out of the 50 units with the best SI. Unit id displayed on top.
@@
Figure A8: Distribution of SI rates for different layers (brain regions rows) over the progress of learning (colors) and change of reward site (purple). Y axes: unit fraction density.
@@
Figure A9: PV correlations across training. Although performance saturates by epochs 5–6, representational changes continue. The last epoch shows the effect of learning a new goal.
Figure A10: Convergence and Remapping. The center of mass (CoM) shift of place fields of individual units from epoch to epoch reveals higher stability in DG and remapping for a new goal.
Figure A11: Occupancy map of LSTM agent’s trajectories across training.
@@
Figure A12: Place fields from the LSTM core (randomly selected units). Each panel shows the normalized spatial activity of one unit.
@@
Figure A13: Population vector correlation across epochs. Color indicates Pearson correlation; rows show epochs 0–4.
@@
Figure A14: Learned dense input features in LSTM agent. Each tile shows the spatial receptive field of one feature.
Figure A15: Displacement-dependent Pearson correlation kernel of population vectors. Axes show spatial shifts $(\Delta x,\Delta y)$; values are averaged correlations.
@@
Figure A16: Training performance with different environment action repeats and number of policies in population based training. Left: environment frame skip / action repeat = 8. Right: environment frame skip / action repeat = 4. The frame skip controls the fine-graininess of actions, which is not required for the current navigation task. On the other hand, larger frame skip effectively make sequences propagate over longer traversals.